OpenStudy (anonymous):

use Laplace transforms to solve the initial value problem: y''+x+2y=0 x''+2x+4y=0 x(0)=y(0)=1 x'(0)=y'(0)=0

OpenStudy (anonymous):

hello again and thanks again

OpenStudy (anonymous):

would you have to start by setting the two equations equal to eachother or solve for one then substitute into the other?

OpenStudy (accessdenied):

Hmm, I haven't seen one like this before. :P I am thinking we should first take the Laplace transform of each equation and get it down to the Laplace transforms of x and y (taking care of those derivatives again). We would get something like a two-variable system of linear equations, I believe.

OpenStudy (accessdenied):

\( y" + x + 2y = 0 \) \( \mathcal{L} (y" + x + 2y) = 0 \) \( \mathcal{L}(y") + \mathcal{L}(x) + 2 \mathcal{L}(y) = 0 \) \( s^2 \mathcal{L}(y) - s y(0) - y'(0) + \mathcal{L}(x) + 2 \mathcal{L}(y) = 0 \) And similarly for the other equation as well.

OpenStudy (accessdenied):

\( x" + 2x + 4y = 0 \) \( \mathcal{L} (x" + 2x + 4y) = 0 \) \( \mathcal{L}(x") + 2 \mathcal{L}(x) + 4 \mathcal{L}(y) = 0 \) \( s^2 \mathcal{L}(x) - s x(0) - x'(0) + 2 \mathcal{L}(x) + 4 \mathcal{L}(y) = 0 \)

OpenStudy (accessdenied):

#1: \( s^2 \mathcal{L}(y) - s \times 1 - 0 + \mathcal{L}(x) + 2 \mathcal{L}(y) = 0 \) \( (s^2 + 1) \mathcal{L}(y) + \mathcal{L}(x) - s = 0 \) #2: \( s^2 \mathcal{L}(x) - s \times 1 - 0 + 2 \mathcal{L}(x) + 4 \mathcal{L}(y) = 0 \) \( 4 \mathcal{L}(y) + (s^2 + 2) \mathcal{L}(x) - s = 0 \) If we just casually treat s as a constant, we just have \(\mathcal{L}(x) \) and \( \mathcal{L}(y) \) as the 'variables' to solve for.

OpenStudy (anonymous):

in the first one, isn't it s^2+2?

OpenStudy (accessdenied):

Oh, yes, sorry. \( (s^2 + 2) \mathcal{L}(y) + \mathcal{L}(x) - s = 0 \) \( 4 \mathcal{L}(y) + (s^2 + 2) \mathcal{L}(x) - s = 0 \) Thank you. I'm still thinking about this method. :P

OpenStudy (accessdenied):

Multiply by (s^2 + 2) so that we can eliminate the \(\mathcal{L}(x) \) \( (s^2 + 2) \left( (s^2 + 2) \mathcal{L}(y) + \mathcal{L}(x) - s \right) = 0 \) \( (s^2 + 2)^2 \mathcal{L}(y) + (s^2 + 2) \mathcal{L}(x) - s (s^2 + 2) = 0 \) \( (s^2 + 2)^2 \mathcal{L}(y) + (s^2 + 2) \mathcal{L}(x) - s (s^2 + 2) = 0 \) \( 4 \mathcal{L}(y) + (s^2 + 2) \mathcal{L}(x) - s = 0 \) Subtract the equations to eliminate \( ((s^2 + 2)^2 - 4) \mathcal{L}(y) - s(s^2 + 2) + s = 0 \) \( ((s^2 + 2)^2 - 4) \mathcal{L}(y) - s(s^2 + 2 - 1) = 0 \) \( ((s^2 + 2)^2 - 4) \mathcal{L}(y) - s(s^2 + 1) = 0 \) \( ((s^2 + 2)^2 - 4) \mathcal{L}(y) = s(s^2 + 1) \) \( \displaystyle \mathcal{L}(y) = \frac{s(s^2 + 1)}{(s^2 + 2)^2 - 4} \) Very interesting...

OpenStudy (accessdenied):

So, that definitely looks like we can break it down a bit because of the denominator being a difference of squares and thus factoring a bit. We can also use that to find \(\mathcal{L}(x) \) if we substitute into one of the equations. Does that make sense?

OpenStudy (anonymous):

I think so. It looks like an interesting equation to try to break down. do you know of anything we can use off the top of your head? I'm thinking of a couple but seem to not be able to completely reduce it.

OpenStudy (accessdenied):

Well, since we'll be using inverse Laplace transforms on it, the technique we used before should be helpful -- Partial fractions. First, though, we need to get the actual factors in the denominator. \( \displaystyle \mathcal{L}(y) = \frac{s(s^2 + 1)}{(s^2 + 2)^2 - 2^2} \) \( \displaystyle \mathcal{L}(y) = \frac{s(s^2 + 1)}{(s^2 + 2 + 2)(s^2 + 2 - 2)} \) \( \displaystyle \mathcal{L}(y) = \frac{s(s^2 + 1)}{(s^2 + 4)s^2} \) Note that we can cancel a bit now. \( \displaystyle \mathcal{L}(y) = \frac{s^2 + 1}{(s^2 + 4)s} \)

OpenStudy (accessdenied):

Page glitch. o.o I see that if we distribute the denominator to each individual part, we can cancel in s^2/((s^2+4)s) to get an identity for inverse Laplace, and then use partial fractions on the remaining 1/((s^2+ 4)s). Can you see how that will work out?

OpenStudy (accessdenied):

\( \displaystyle \mathcal{L}(y) = \frac{s^2 + 1}{(s^2 + 4)s} \) \( \displaystyle \mathcal{L}(y) = \frac{s^2}{(s^2 + 4) s} + \frac{1}{(s^2 + 4) s} \) \( \displaystyle \mathcal{L}(y) = \frac{s^\cancel{2}}{(s^2 + 4) \cancel{s}} + \frac{1}{(s^2 + 4)s} \) \( \displaystyle \mathcal{L}(y) = \frac{s}{s^2 + 4} + \frac{1}{(s^2 + 4)s} \) Like that. :)

OpenStudy (anonymous):

so the first one is cos(2t) but what's the second? I can't seem to find a similar one in the table that I have

OpenStudy (accessdenied):

We need to do a bit more to that one. Notice that we have two factors in the denominator, so now we have to use Partial fractions. Do you recall that method we used earlier on (2s^2 + 7s + 45)/(s(s^2 + 2s + 5)) ? We set it equal to A/s + (Bs + C)/(s^2 + 2s + 5). The same thing will be done here. \( \displaystyle \frac{1}{(s^2 + 4)s} = \frac{A}{s} + \frac{Bs + C}{s^2 + 4} \)

OpenStudy (anonymous):

sorry I'm still writing a few things down. I'll be caught up in a second :)

OpenStudy (accessdenied):

Alright. :)

OpenStudy (accessdenied):

From here, it is really the same process as last time. We just have to pretend the left-hand-side has 0s^2 + 0s + 1 when we equate the coefficients.

OpenStudy (anonymous):

so A=1/4, B=-1/4, and C=0 is what I got

OpenStudy (accessdenied):

Yep, that is what I have as well. :) So, we can then split it up like so: \( \displaystyle \frac{1}{4} \frac{1}{s} + \frac{1/4\; s + 0}{s^2 + 4} \) Oh look, we have another single-unit denominator. That looks like a couple of our transforms now. :)

OpenStudy (accessdenied):

and I forgot a negative there, it should have been there at the 1/4s. :P

OpenStudy (accessdenied):

\( \displaystyle \mathcal{L}(y) = \frac{s}{s^2 + 4} + \frac{1}{(s^2 + 4)s} \) \( \displaystyle \mathcal{L}(y) = \frac{s}{s^2 + 4} + \frac{1}{4} \frac{1}{s} - \frac{1}{4} \frac{s}{s^2 + 4} \) Looks good now. :)

OpenStudy (anonymous):

so y=cos(2t)+1/4-1/4cos(2t), right? So how would we go about solving for x?

OpenStudy (accessdenied):

you may like to combine the cos(2t) and -1/4 cos(2t) in that part into 3/4 cos(2t). Anyways, we need to take the Laplace transform of y and basically substitute it back into one of the two equations we had: \( (s^2 + 2) \mathcal{L}(y) + \mathcal{L}(x) - s = 0 \) \( 4 \mathcal{L}(y) + (s^2 + 2) \mathcal{L}(x) - s = 0 \) We can use any form of the Laplace transform of y in the equation of choice. I am thinking we should use the second equation. I feel like we should use \( \displaystyle \mathcal{L}(y) = \frac{s^2 + 1}{(s^2 + 4)s} \) because its more compact for substituting. \( \displaystyle 4 \color{#aaaa55}{\frac{s^2 + 1}{(s^2 + 4)s}} + (s^2 + 2) \mathcal{L}(x) - s = 0 \)

OpenStudy (anonymous):

wow... that looks like a mess

OpenStudy (accessdenied):

Actually, I am not sure if the expanded form might be easier or not... I could check, I suppose.

OpenStudy (accessdenied):

Actually, if we simplify the expanded form a bit, it should work out better. Sorry to backtrack... >.> \( \displaystyle \mathcal{L}(y) = \frac{s}{s^2 + 4} + \frac{1}{4} \frac{1}{s} - \frac{1}{4} \frac{s}{s^2 + 4} \) \( \displaystyle \mathcal{L}(y) = \frac{3}{4} \frac{s}{s^2 + 4} + \frac{1}{4} \frac{1}{s} \) Because then we have: \( \displaystyle 4 \left( \frac{3}{4} \frac{s}{s^2 + 4} + \frac{1}{4} \frac{1}{s} \right) + (s^2 + 2) \mathcal{L}(x) - s = 0 \) \( \displaystyle 3 \frac{s}{s^2 + 4} + \frac{1}{s} + (s^2 + 2) \mathcal{L}(x) - s = 0 \) I mean, it will work out either way but it would be nice to have the least issues while working. Either way, we got some work to do with dividing off the (s^2 + 2) on the Laplace transform.

OpenStudy (accessdenied):

\( \displaystyle (s^2 + 2) \mathcal{L}(x) = s - \frac{1}{s} - 3 \frac{s}{s^2 + 4} \) \( \displaystyle \mathcal{L}(x) = \frac{s}{s^2 + 2} - \frac{1}{s(s^2 + 2)} - 3 \frac{s}{(s^2 + 4)(s^2 + 2)} \) We pretty much have to use partial fractions on the latter two terms to break those up into useful forms now. The first one is virtually a repeat of the one we did previously, but the second one is just slightly more difficult because there are two irreducible quadratics in the denominator.

OpenStudy (anonymous):

what would the very first one be? would we use a square root of 2 for that one?

OpenStudy (accessdenied):

Yeah. Since it is an s^2 + (sqrt(2))^2, we need that sqrt2.

OpenStudy (anonymous):

@AccessDenied A = 1/2, B=-1/2, and C=0? So we have the first and the second. I'm drawing a blank on the third one. Would the top part be as+B and Cs+D?

OpenStudy (accessdenied):

Yes, the first coefficients are correct. Also, yes, we have (As + B)/... + (Cs+D)/...

OpenStudy (accessdenied):

I apologize in advance if something happens and I disconnect soon. I seem to be disconnecting from OS every couple minutes. o.O

OpenStudy (accessdenied):

\( \displaystyle \frac{s}{(s^2+4)(s^2+2)} = \frac{As + B}{s^2+4} + \frac{Cs + D}{s^2+2} \) The set up You will have a four-variable system of equations, but luckily again it does not seem to be too difficult to solve. It should turn out to be almost like two two-variable systems.

OpenStudy (anonymous):

I got A=-1/2,B=0, C=1/2, and D=0.

OpenStudy (accessdenied):

Yep, that looks correct to me. :) Now we substitute: \( \displaystyle \mathcal{L}(x) = \frac{s}{s^2 + 2} - \left( \frac{1}{2} \frac{1}{s} - \frac{1}{2} \frac{s}{s^2+2} \right) - 3 \left( \frac{1}{2} \frac{s}{s^2 + 2} - \frac{1}{2} \frac{s}{s^2 + 4} \right) \) \( \displaystyle \mathcal{L}(x) = \frac{s}{s^2 + 2} - \frac{1}{2} \frac{1}{s} + \frac{1}{2} \frac{s}{s^2+2} - 3 \frac{1}{2} \frac{s}{s^2 + 2} + 3 \frac{1}{2} \frac{s}{s^2 + 4} \)

OpenStudy (anonymous):

I get x=cos(sqrt2 t)-1/2+1/2 cos(sqrt2 t)+3/2 cos(2t)-3/2 cos(sqrt2 t)

OpenStudy (accessdenied):

Hmm, you might be able to simplify that up quite a bit.

OpenStudy (anonymous):

wow... ya you're right. not sure why I didn't...

OpenStudy (anonymous):

x=3/2 cos(2t) - 1/2, yes?

OpenStudy (accessdenied):

Yes! I actually did the simplifications before taking the inverse, which also works out... \( \displaystyle \mathcal{L}(x) = \frac{s}{s^2 + 2} - \frac{1}{2} \frac{1}{s} + \color{#aa5555}{\frac{1}{2} \frac{s}{s^2+2} - \frac{3}{2} \frac{s}{s^2 + 2}} + 3 \frac{1}{2} \frac{s}{s^2 + 4} \) 1/2 - 3/2 = -1 \( \displaystyle \mathcal{L}(x) = \frac{s}{s^2 + 2} - \frac{1}{2} \frac{1}{s} - \frac{s}{s^2 + 2} + 3 \frac{1}{2} \frac{s}{s^2 + 4} \) \( \displaystyle \mathcal{L}(x) = \cancel{\frac{s}{s^2 + 2}} - \frac{1}{2} \frac{1}{s} - \cancel{\frac{s}{s^2 + 2}} + 3 \frac{1}{2} \frac{s}{s^2 + 4} \) Incredible cancelling! \( \displaystyle \mathcal{L}(x) = -\frac{1}{2} \frac{1}{s} + 3 \frac{1}{2} \frac{s}{s^2 + 4} \)

OpenStudy (anonymous):

ya definitely. Thank you very much and do you have some more time to help or do you have something else to do? If you have some more time, I have two more questions, if not, that's fine too.

OpenStudy (accessdenied):

Well, I might take a break for a while. Are they similar to these last two where you may try those on your own and we can check the work?

OpenStudy (anonymous):

the first one has to do with the translation(or shifting)theorem and the second has to do with the differential of transforms theorem. I'm not too familiar with either of them.

OpenStudy (accessdenied):

Alright. Well, I am probably going to take around a 10 to 20 minute break and then I will check back again. I think at this time there may be more people around who are keen on DE, though, so you could post the question and see if somebody else will help you as well. :)

OpenStudy (anonymous):

will do. Maybe I'll see you again. Thanks a lot for the help though :)

OpenStudy (accessdenied):

You're welcome! Thanks for the interesting problems here. :D

OpenStudy (anonymous):

I've got more if I see you again!